UT Feb 04 2021



Night Partition

D/E Time - twilight until 21hr - handover to PEPSI (local)

PEPSI 21hr - 03:30hr (local)

D/E Time - 03:30 to twilght (local)

Script Location: /home/lbto/Service_Observing/D_E_Feb_Mar_2021/

Night Plan

  • Configure for Mixed-Mode MODS-1 SX - LBC-Red DX
  • MODS1+LBCR mixed-mode on-sky test (OPK)
    • 0.5-1 hr
    • Run a long MODS observation with MOS-type alignment (i.e., dtheta, dx, dy) and a simultaneous LBCR dithered sequence.
      • Make sure we can stop LBCR, refocus/recollimate, and then load & play a new LBCR OB.
    • priority=high (for an LBTB program), seeing=poor, clouds=don't care
  • Reconfigure for Bino-MODS
  • MODS1 Active Optics parameter change (JMH) - 1 hr
    • MODS1 collimating with ~3 different guide stars (don't care what MODS is looking at)
    • priority=medium-low, seeing=average, clouds=don't care
  • UGC 4079 - MODS Sci Program - filler to reach PEPSI handover time (BR) - 1hr-ish
  • Configure for PEPSI until 03:30
  • Reconfigure at 03:30 (local) to LBC-Bino
  • LBC Pointing Hysteresis and Settling Test (JMH/OPK)
    • needs 1 hour max in any conditions where we are open
    • priority=medium-low (interesting for TMS)
  • 2020 SO LBC observations (CV) - 1hr

Pre-Open

  • Both LBCs are run up. Right side authorized for LBC and LBCR connected. Left side is authorized for MODS1. LBC IRAF, IDL and RB_Science are being run from robs, but I have the scripts on obs3 and so LBC UI on obs3.
  • I made a "custom" 9pt dither with 20-30" gaps for the initial and repeat scripts.


Night Log

01:45 The enclosure remains closed, due to clouds. Earlier there were high winds, but these have calmed down a bit.

02:12 No, winds reached 23.3 m/s just a while ago... still closed.

02:43 Wind gust of 22.3 m/s.

03:56 Time to start reconfiguring to PEPSI for the time-critical observation. We remain closed due to high winds. Heejoo turned the TMS laser off.

08:18 - still closed- winds 31m/s+, humidity hit 99% now at 79%. Still clouds on allsky

09:59 - I.Ilyn handed over from PETS/PEPSI to D/E. Now reconfiguring for LBCs. Greg verified that the elevation limit was "automagically" changed to 30 deg.

11:34 - Wind gusts are still very high. Three times already in the past 1.5 hours, both Barry's and my x2go sessions to the mountain and the polycom have all dropped out.

11:37 - And again, the x2go sessions & polycom dropped out.

-TMS laser system is on - still have 35m/s+ winds

12:43 Closed Dome test of sending a preset to a star in an "OSA" catalog created by pngui.
  • Create the catalog
    • ssh telescope@obs1
    • pngui
      • Click CREATE tab
      • Enter coordinates where you want to slew, then
      • Click "Create Catalog Entry" and "SAVE". This will search for stars within "Err" of the input coordinates and create a catalog entry "OSA###.catalog" under /lbt/data/catalogs/lbto/Scratch/OSA/catalogs
  • To Slew to a star in the new catalog:
    • open 2 IRAF sessions and set up LBTtools:
      • camera = LBC (or RedLBC or BlueLBC)
      • Telside = "Left" in one and "Right" in the other IRAF session
    • load LBTtools and Observe
    • epar newpoint to enter the catalog name, TelMode and gstarID
    • Issue the command, newpoint BC=1, on both sides.

-- BarryRothberg - 03 Feb 2021
Topic revision: r9 - 04 Feb 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback